Blue Angel, Inc., a private firm in the holiday gift industry, is considering a new project. The company currently has a target debt–equity ratio of .45, but the industry target debt–equity ratio is .40. The industry average beta is 1.20. The market risk premium is 8 percent, and the risk-free rate is 6 percent. Assume all companies in this industry can issue debt at the risk-free rate. The corporate tax rate is 40 percent. The project requires an initial outlay of $680,000 and is expected to result in a $100,000 cash inflow at the end of the first year. The project will be financed at the company’s target debt–equity ratio. Annual cash flows from the project will grow at a constant rate of 6 percent until the end of the fifth year and remain constant forever thereafter.

Calculate the NPV of the project.
Answer #1

Initial outlay = 3680000 Company Debt: equity = 0.45 Let west ben savity = 1-X x = vous 1-n x=0.45 0.45X lousna oous ! na 0.3Edination Year o ก of Cashflauf Coosh flow (6X0,000) 100000 100000 * 0,06) : 100000 (1.00)? ? 1000000 m (1.06)3 2 100000 (1.0

Please give good rating

comment if u have any query ..